### v2

# Enabling Large Language Models to Perform Power System Simulations with Previously Unseen Tools: A Case of Daline

Abstract
The integration of experiment technologies with large language models (LLMs) is transforming scientific research, offering AI capabilities beyond specialized problem-solving to becoming research assistants for human scientists. In power systems, simulations are essential for research. However, LLMs face significant challenges in power system simulations due to limited pre-existing knowledge and the complexity of power grids. To address this issue, this work proposes a modular framework that integrates expertise from both the power system and LLM domains. This framework enhances LLMs’ ability to perform power system simulations on previously unseen tools. Validated using 34 simulation tasks in Daline, a (optimal) power flow simulation and linearization toolbox not yet exposed to LLMs, the proposed framework improved GPT-4o’s simulation coding accuracy from 0% to 96.07%, also outperforming the ChatGPT-4o web interface’s 33.8% accuracy (with the entire knowledge base uploaded). These results highlight the potential of LLMs as research assistants in power systems.
Index Terms:
Power Systems, Large Language Models, Simulation, Co-researcher, Daline
I Introduction
Combining laboratory automation technologies with large language models (LLMs) enables automated execution of scientific experiments [1]. Related advances span the fields of mathematics, chemistry, and clinical research, including mathematical algorithm evolution [2], geometry theorem proving [3], chemical experiment design and execution [1], as well as the development and validation of machine learning approaches for clinical studies [4]. These recent achievements signal a new research paradigm, positioning AI as a research assistant for humans with natural language communication abilities, rather than merely a specialized problem solver as in the past.
Establishing LLMs as research assistants has significant potential for advancing power system studies, which heavily rely on simulations. To develop LLM-based assistants for power systems, it is crucial to equip LLMs with the ability to perform these simulations, a capability not inherent to LLMs. For instance, even GPT-4 often struggles to create small distribution grids using OpenDSS [5] or writing code for simple power flow problems [6]. This limitation is evident despite the widely available knowledge on optimal power flow problems. However, existing studies mainly focus on conceptualizing [7], demonstrating [8, 7], and evaluating [5, 6] LLMs’ capabilities in generating power system simulation codes, rather than systematically developing and enhancing their ability to perform these simulations.
To bridge this gap and resolve the above limitation of LLMs, this letter first argues that establishing simulation capabilities in LLMs requires a modular framework that integrates and coordinates multiple techniques. Beyond explicit elements like (i) prompt engineering to enhance LLM performance [8] and (ii) retrieval-augmented generation (RAG) to incorporate specific power systems knowledge into LLMs [9, 7, 6], this framework should also consider often overlooked implicit factors: (iii) the supplementary design of the simulation toolbox (including automated syntax checking and error reporting, and the architecture of the tool’s knowledge base), and (iv) the natural language interactive feedback loop between LLMs and the simulation executor.
Building on this concept, this letter proposes a four-module framework to enable LLMs to perform power systems simulations using a simulation toolkit not previously exposed to LLMs. This framework integrates specialization from both the power system and LLM domains. Subsequently, the proposed framework, though being toolbox-independent, is applied to the Daline [^1] Centered on power system simulations, Daline includes functionalities such as (optimal) power flow data generation, data pollution, data cleaning, data normalization, method selection, method customization, model linearization, model evaluation, and result visualization. It supports a large amount of standard power system cases, 57 power flow linearization methods, and over 300 customizable options. See https://www.shuo.science/daline for more details. toolbox [10] for validation, as Daline was released after the latest updates of any LLMs tested in this letter. Results show that the proposed framework significantly enhances the simulation performance of LLMs. This improvement is a cumulative effect of incorporating multiple techniques, as presented in the following.
II Proposed Modular Framework
Figure 1: Proposed framework with techniques indexed from 1 to 10. $N$ is the number of feedback iterations and $N_{max}$ is the maximum number of iterations.
The proposed framework consists of four modules with multiple techniques: (i) prompt design, (ii) enhanced RAG design, (iii) LLM-oriented simulation toolbox supplementary design, and (iv) feedback loop design, as illustrated in Fig. 1 and detailed below.
II-A Prompt Design
To support the LLM to understand its role and purpose, we customized several prompt engineering techniques, including chain of thought prompting [11] and few-shot prompting [12], for toolbox-based simulations. Beyond clarifying the LLM’s role and primary functionality, we defined its actions step-by-step: (i) identifying simulation functions, (ii) syntax learning, (iii) locating necessary parameters/options, (iv) writing code, (v) providing references, and (vi) making conclusions, with examples for clarity. The major syntax of the toolbox is also explained in the role prompting. While the above prompt engineering techniques mainly come from the LLM domain, the design of actions, specifics of each prompt, and customization of examples heavily depend on expertise in the power system simulation tool. For the complete prompt, see the supplementary file “role_description.txt.”
II-B Enhanced RAG Design
For power system simulation tools unfamiliar to LLMs, it is necessary to impart specific knowledge about the tool. RAG [9], a cost-effective approach, can integrate this information into LLMs while reducing hallucinations. Existing studies have used the standard RAG (powered by LangChain) for long-context question answering [6] and non-specific code generation [7] in power systems. The standard RAG procedure includes external knowledge chunking (splitting external documents into smaller pieces), text embedding (converting texts into vectors using existing text2vec neural networks [^2] The text2vec model we used in this study is from here.), and information retrieval (finding information in the vector space that matches the whole user request) [6]. However, user requests often involve multiple functions and parameters spread across documents. Simply using the whole request sentence for retrieval may not collect enough semantic information across different sources and granularities.
User requests for simulations typically include two critical elements: the functions to be used and the parameters to be set. Hence, to address the above issue, we developed a prompt-based query planning strategy for LLMs. First, we enable LLMs to decompose long requests into sub-requests, each corresponding to a specific simulation function or parameter. Then, we enable LLMs to map each sub-request to a keyword representing the related function or parameter for parallel retrieval. This strategy, leveraging the synergy between LLM and power system simulation expertise, is integrated into the standard RAG structure, resulting in an enhanced RAG architecture that improves the retrieval of critical information from multiple knowledge sources, as shown in Fig. 1. The complete query-planning prompt is provided in the supplementary file “query_planning.txt”.
II-C LLM-oriented Simulation Toolbox Supplementary Design
In addition to the previously presented designs, hundreds of tests in our study show that supplementary design for the simulation toolbox is also needed to reliably enable LLMs to perform power system simulations. This includes (i) developing a RAG-friendly knowledge base, and (ii) a syntax checking and error reporting system, both for the toolbox.
Specifically, power system simulation toolboxes usually have user manuals detailing functions, parameters, syntax, and examples. While can be used as external knowledge for RAG, user manuals are designed for human readability and often spread critical information across different pages, tables, and figures, making them unsuitable for information retrieval. Hence, we propose adding two RAG-friendly documents: one lists all supported parameters/options in the toolbox. Each is written in one line, with its name, default value, explanation, and associated functions/methods (acts as a locator to help RAG link parameters with functions/methods). Another contains all code examples from the manual, organized in a clear structure. These documents help RAG capture more precise information than the user manual alone. All documents are available in the supplementary file “knowledge_base.zip.”
In addition, toolboxes should pre-check syntax and input formats of each function before deeper code execution. Common syntax errors can be corrected internally, while other errors should provide clear messages about the original cause and troubleshooting hints. Although some toolboxes may already have such features, extra attention and further effort are needed when the users are LLMs rather than humans. These features, combined with the feedback loop discussed below, aid LLMs in reasoning and correcting their coding errors automatically.
II-D Feedback Loop Design
LLMs can make mistakes, but a feedback loop between the simulation executor and LLMs can iteratively correct them. With an established syntax checking and error reporting system, the feedback design amounts to providing a comprehensive error report to LLMs, including (i) the problematic code, (ii) a clear error message, (iii) troubleshooting hints, (iv) a request to correct the code, (v) reminders of common mistakes, and (vi) an organized chat history. This feedback design significantly improves the success rate of LLMs with weaker comprehension abilities, such as GPT-3.5.
III Case Study
To verify the proposed framework, 20 schemes (technique indices correspond to Fig. 1) listed in Table I were evaluated using 34 power system simulation tasks in Daline. These tasks, including 27 normal and 7 complex requests written in natural language, cover the full functionality of Daline, from generating AC power flow datasets to data pollution, cleaning, normalization, and power flow linearization. Complex requests also compare and rank the accuracy and computational efficiency of various methods with different settings for training, testing, and visualizing. Each task was tested independently. The complete set of task requests is provided in the supplementary file “simulation_request.txt,” and experiment records are in “result.zip.” Code will be open-source upon acceptance.
For performance evaluation, each scheme has 3 attempts ($N_{max}=3$) per simulation request. A scheme earns 1 point per attempt for exact correct code without irrelevant settings, 0.5 points for correct code with irrelevant settings, and 0 points for code with mistakes. Subsequent attempts are made only if the previous one encounters execution errors. Attempts not triggered get the same score as the last attempt. Coding accuracy per scheme is defined as the total points earned divided by the possible highest score ($34\times 3=102$ here), resulting in an accuracy level between 0% and 100% per scheme.
Figure 2: Simulation Coding Accuracy Comparison of Evaluated Schemes (the feedback loop is enabled for all schemes).
TABLE I: Evaluated Schemes (Technique Index Numbers From Fig. 1)
Scheme Techniques LLM RAG
GPT-4o-Full 1,2,3,5,6,7,9,10 GPT-4o (API) Proposed
GPT-3.5-Full 1,2,3,5,6,7,9,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NRPL 1,2,3,5,6,8,9,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NRM 1,2,3,5,7,9,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NG 1,2,3,5,6,7,9 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NK 1,2,3,4,6,7,9,10 GPT-3.5-Turbo (API) Standard
GPT-3.5-NC 2,3,5,6,7,9,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NKC 2,3,4,6,7,9,10 GPT-3.5-Turbo (API) Standard
GPT-3.5-NRE 1,2,3,5,6,7,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NRP 1,2,3,5,6,9,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NREP 1,2,3,5,6,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NS 1,3,5,6,7,9,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-Prompt 1,2,3,10 GPT-3.5-Turbo (API) -
GPT-3.5-NCS 3,5,6,7,9,10 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NGS 1,3,5,6,7,9 GPT-3.5-Turbo (API) Proposed
GPT-3.5-NKS 1,3,4,6,7,9,10 GPT-3.5-Turbo (API) Standard
ChatGPT-4o-R 6,7,9,10 ChatGPT-4o Web Interface OpenAI
GPT-4o-R 4,6,7,9 GPT-4o (API) Standard
GPT-4o-Sole 1,10 GPT-4o (API) -
GPT-3.5-Sole 1,10 GPT-3.5-Turbo (API) -
The accuracy performance of the evaluated schemes is shown in Fig. 2. Both GPT-3.5-Sole and GPT-4o-Sole have zero accuracy, indicating they have not encountered Daline before. GPT-4o-R achieves only 12.25%, suggesting that only using the standard RAG [7, 6] is unreliable for LLMs in power system simulations. Even with OpenAI’s official RAG tool and the entire knowledge base, ChatGPT-4o-R’s accuracy is only 33.82%. However, with the proposed framework, GPT-4o-Full achieves 96.07% accuracy. Importantly, the bold black polyline in Fig. 2 shows that incorporating more techniques from the proposed framework significantly improves LLMs’ performance.
Additionally, Fig. 2 also highlights the impact of individual techniques on accuracy. For example, the enhanced RAG structure raises accuracy from 74.01% (GPT-3.5-NK) to 81.37% (GPT-3.5-Full). Without few-shot prompting, accuracy improves from 20.58% (GPT-3.5-NKS) to 45.09% (GPT-3.5-NS) after using the enhanced RAG structure. Once few-shot prompting is implemented, accuracy jumps from 45.09% (GPT-3.5-NS) to 81.37% (GPT-3.5-Full). Furthermore, only using RAG-friendly documents as the knowledge base enhances performance (75.49% accuracy for GPT-3.5-NRM) compared to only using the user manual (60.29% accuracy for GPT-3.5-NREP). Similarly, syntax error checking and the reporting system combined with few-shot prompting yield significant improvements, as shown by the gray polyline in Fig. 2. Overall, the accuracy ranking (GPT-3.5-Full $>$ GPT-3.5-NRPL $>$ GPT-3.5-NRM $>$ GPT-3.5-NG $>$ GPT-3.5-NK $>$ GPT-3.5-NC $>$ GPT-3.5-NRE $>$ GPT-3.5-NRP $>$ GPT-3.5-NS) summarizes the contributions of individual techniques. This also demonstrates that achieving high accuracy is a cumulative result of multiple techniques, emphasizing the necessity of a systematic framework with various techniques to enable LLMs to reliably perform complex power system simulations.
IV Conclusion
This letter proposes a modular framework to enable LLMs to perform power system simulations on previously unseen tools. The framework includes four modules with multiple techniques. Evaluated across 34 diverse tasks in the Daline toolbox, the framework increased coding accuracy for GPT-4o from 0% to 96.07%, surpassing the ChatGPT-4o web interface’s 33.82% accuracy. The impacts of individual techniques have been quantified using 20 different combinations of LLM versions and proposed techniques, demonstrating that high accuracy is achieved through the cumulative effect of multiple techniques. This underscores the necessity of a systematic framework with various techniques to enable LLMs to perform complex power system simulations reliably. Overall, this work highlights the potential for LLMs as research assistants in power systems. Since the proposed framework is currently limited to using a single simulation toolbox, future research will focus on generalizing the framework to accommodate multiple power system simulation tools.